Fermi National Laboratory


Accelerator Update

Monday May 10
The day shift began with Main Injector (MI) off due to power supply problems. Tevatron (Tev) is waiting to go into shot setup.

At 10:23 AM CDF and D0 both used this downtime to make a 1 hour controlled access.

MI was back on and in a stable state at 1:04 PM.

Shot setup for the Tev began at 1:40 PM.

At 3:31 PM Operations established store #3481 with an initial luminosity of 59.20 E30.

Antiproton Source (Pbar) returns to stacking, beam to MiniBoone at 4:17 PM.

Beam to MCenter at 6:03 PM.

Tuesday May 11
The Midnight shift began with store #3481 colliding, stacking, and with beam to MiniBoone and MCenter.

Turned on beam to MTest at 9:20 AM.

At 12:41 PM the Pbar stack went from 145ma to 58ma while going to the reverse proton mode. No obvious cause was found for the Pbar loss.

MCenter turned off at 1:28 PM.

Mtest turned off at 6:29 PM.

Wednesday May 12
The Midnight shift began with store #3481 colliding, stacking, and with beam to MiniBoone and MCenter.

M01RAW tripped off at 2:38 AM. The AC breaker was reset and beam resumed to MCenter. This device continued to trip every hour or so through out the shift.

Store # 3481 was terminated at 5:31 AM.

Today's Plans
Tev studies during the day, followed by shot setup this evening. M01RAW will be repaired.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/12/2004   email Fermilab